OpenStudy (anonymous):

Find the equation of a circle whose center is at the origin and radius is 4.

OpenStudy (jmark):

Equation of Circle: (x - h)^2 + (y - k)^2 = r^2 where, [(h,k) is the center of the circle at the origin] r is the radius, (h,k) = (0,0), r = 4 (x - 0)^2 + (y - 0)^2 = 4^2 x^2+y^2 = 16 the equation of the circle x2+y2 = 16 ICSE Results 2014 - http://goo.gl/aZye9y